# Defining implication relation for classical logic

### v11

@@ 10 $\vDash\chi[\phi\mapsto\psi]$ 9: Definition
Suppose a conditional proof $\phi\vdash\psi$ of IRL satisfies that if a member of its proof sequence is a theorem then it is a tautology, then $\phi\to\psi$ is a tautology. That is, the rule of conditional proof is tautologousness-preserving.
@@
Only the three primitive rules need considered in the meta proof: Conjunction introduction $\{\phi,\psi\}\Rightarrow\phi\wedge\psi$, Conjunction elimination $\phi\wedge\psi\Rightarrow\phi$, Replacement property $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$.
@@
    • Case 1: $\psi$ is a theorem. In this case, $\psi$ is a tautology, so $\phi\to\psi$ must be a tautology as shown below:
@@ 7     $\vDash\phi\to\chi$ 6: Definition
    9     $\vDash\phi\to\chi$ 8: Induction hypothesis
@@ 10 $\vDash\phi\to\chi$ 2,3–7,8–9: Proof by exhaustion
    This result is used in the following five cases of $\psi$.
@@ 11 $\vDash\phi\to\psi$ 10: Definition
    Thus, for all $n\geq 1$, suppose that for all $k=1,...,n$, any conditional proof $\phi\vdash\psi$ of length $k$ satisfying the “theorem-is-tautology” condition results in tautological $\phi\to\psi$, then any length-$(n+1)$ conditional proof $\phi\vdash\psi$ leads to tautological $\phi\to\psi$ as well, under the same condition.
@@
From Meta Theorem 4.4.1, Meta Theorem 4.4.2 and Meta Theorem 4.4.3, it can be concluded that IRL is sound with regard to its semantics specified by IRLA.
@@
    All axioms of Boolean lattice have their counterparts in IRL. For example, IRLA axiom $x\wedge y=y\wedge x$ corresponds to IRL axiom $\phi\wedge\phi\leftrightarrow\psi\wedge\phi$.
@@
    The IRLA axiom $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$ (one of the two conditions for defining $\leftrightarrow$ in IRLA) corresponds to IRL theorem $(\phi\leftrightarrow\psi)\wedge\chi\to\chi[\phi\mapsto\psi]$ (from primitive rule $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$).
@@
    The IRLA axiom $x\to y=x\wedge y\leftrightarrow x$ (defining axiom for $\to$) corresponds to IRL axiom $(\phi\to\psi)\leftrightarrow(\phi\wedge\psi\leftrightarrow\phi)$.
@@
For all the primitive rules of IRLA, i.e. $x=y\Rightarrow x\leftrightarrow y=1$ and $\{s=t,u=v\}\Rightarrow(u=v)[s\mapsto t]$, if their premises have corresponding theorems in IRL, so are their conclusions.
@@
    $x=y\Rightarrow x\leftrightarrow y=1$: For the premise $x=y$, if $\phi\leftrightarrow\psi$ is a theorem, then $(\phi\leftrightarrow\psi)\leftrightarrow\top$ is a theorem by the rule of theorem representation in IRL, and the latter is just the counterpart in IRL for the conclusion $x\leftrightarrow y=1$.
@@
    $\{s=t,u=v\}\Rightarrow(u=v)[s\mapsto t]$: For the premises $s=t$ and $u=v$, if $\eta\leftrightarrow\zeta$ and $\theta\leftrightarrow\kappa$ are theorems, then $(\theta\leftrightarrow\kappa)[\eta\mapsto\zeta]$, the counterpart in IRL for the conclusion $(u=v)[s\mapsto t]$ in IRLA, must be a theorem as shown below.
    1 $\vdash\eta\leftrightarrow\zeta$ Known
    2 $\vdash\theta\leftrightarrow\kappa$ Known
    3     $\eta\leftrightarrow\zeta$ 1 (assumed theorem)
    4     $\theta\leftrightarrow\kappa$ 2 (assumed theorem)
    5     $(\theta\leftrightarrow\kappa)[\eta\mapsto\zeta]$ 3,4: Replacement
    6 $\vdash(\theta\leftrightarrow\kappa)[\eta\mapsto\zeta]$ 3–5: Unconditional proof
From Meta Theorem 4.4.4 and Meta Theorem 4.4.5, it can be concluded that IRL is complete with regard to its semantics specified by IRLA.
@@ 4.5 Isomorphism between IRL and IRLA
Isomorphism between two systems can be equivalently expressed as they are interpretable in each other in terms of the concept of interpretability (Tarski, 1953).
@@
Since IRL is sound and complete with regard to IRLA, there is a natural isomorphism between them as shown in Table 3.
@@
Formula schema $\eta(\phi,\psi,...)$ $t(x,y,...)$ Term
@@
As IRL is sound and complete, for a formula $\phi$ of IRL, “$\phi$ is valid or is a tautology” is equivalent to “$\phi$ is a theorem’ and “$\phi$ is unsatisfiable or is a contradiction” is equivalent to “$\neg\phi$ is a theorem”; so “$\phi$ is a contingency” is equivalent to “both $\phi$ and $\neg\phi$ are not theorems”.
@@ 4.6 Truth values of implication
Define $x<y$ as $x\leq y$ and $x\neq y$. Then $10<x<$ 1 means that $x$ is (represents) a contingency.
By Theorem 4.3.11, in any cases we have
$0\leq x\to y\leq\neg x\vee y.$
@@
Thus, generally $V_{\phi\to\psi}$ is a subset (that can be an empty set) of $V_{\neg\phi\vee\psi}$. Especially in the case where $\neg\phi\vee\psi$ is contingent, unlike classical logic that always adopts the equality $V_{\phi\to\psi}=V_{\neg\phi\vee\psi}$, $V_{\phi\to\psi}$ can be any subset of $V_{\neg\phi\vee\psi}$, i.e. $V_{\phi\to\psi}\in 2^{V_{\neg\phi\vee\psi}}$, so there are $2^{|V_{\neg\phi\vee\psi}|}$ subsets, of which only one is “correct” for $\phi\to\psi$. The correct $V_{\phi\to\psi}$ is determined by the original meaning of implication – a “certain mechanism” (cf. Subsection 2.2) corresponding to the implication $\phi\to\psi$ considered.
@@
Let $A$ and $B$ are primitive propositions having 4 possible truth-value combinations, i.e. $U=\{\mathrm{TT,TF,FT,FF}\}$. Then $\emptyset\subset V_{\neg A\vee B}=\{\mathrm{TT,FT,FF}\}\subset U$ (so $\neg A\vee B$ is contingent), therefore $V_{A\to B}\in 2^{V_{\neg A\vee B}}=2^{\{TT,FT,FF\}}$, thus there are $82^{|\{\mathrm{TT,FT,FF}\}|}=2^{3}=$ 8 possibilities and only one of them is correct for $V_{A\to B}$.
The next example shows which should be the “correct one” for the T-set of implication $\phi\to\psi$ as a subset of the T-set of $\neg\phi\vee\psi$.
@@ F         F         F         T         F
    T-set of $\neg P\vee Q$ is the set of rows $\{r_{1},r_{2},r_{3},r_{4},r_{6},r_{7}\}$, indicating that it is contingent.
@@
    T-set of $P\to Q$, determined by $O$ rather than $P$ and $Q$, is $\{r_{1},r_{2},r_{3}\}$, a proper subset of that of $\neg P\vee Q$.
In this example, it is the certain mechanism of “in the open air” that corresponds to the inevitable implication relation of “if it rains, then the ground gets wet”. This is a typical example showing that it is not uncommon that $\neg\phi\vee\psi\to(\phi\to\psi)$ is invalid. The example can explain why the classical logic, with its implication unreasonably defined by $(\phi\to\psi)\leftrightarrow\neg\phi\vee\psi$, is doomed to fail in a large number of real situations.
@@ 5.1 Definitions and notations
In the following, view a proof system as the set of its axioms and primitive rules, and use “statement” to mean formula or rule expression of a system; view an object and the singleton of the object the same thing.
@@
Definition 5.1.3 (System extension and restriction).
@@
An extension of $A$ with $X$, denoted as $A+X$, is the system obtained from $A$ by adding each member of $X$ as an axiom or a primitive rule. And $A$ is a restriction of $A+X$ in which $X$ has been removed.
@@
In terms of the above definitions: an extension of a system is a super-system of that system; the deductive equivalent relation “$=$” can be used in a usual way (i.e. it has the properties of reflexivity, symmetry, transitivity and replacement) ; the extension operator “$+$” is commutative and associative. Let $A$, $B$ and $C$ be sets of statements, any of which can be a system, then:
@@ 5.2 Two super-systems of IRLA
Let $BL(X,1,0,\neg,\wedge,\vee,\leq)$ be the Boolean lattice described in Subsection 4.2. $BL0(X,1,0,\neg,\wedge,\vee,\to,\leftrightarrow,\leq)$ and $BL1(X,1,0,\neg,\wedge,\vee,\to,\leftrightarrow,\leq)$ are two extensions of $BL$ defined as
@@
  • BL0 is isomorphic to a “contra-classical” logic, denoted as CL0, obtained by translating BL0 to a logic in terms of Table 3 so that $BL0\cong CL0$. The logic CL0 is contra-classical because the definitions of implication $x\to y=\mu(x,y)$ in BL0 and $x\to y=\neg x\vee y$ in BL1 are contradictory in the general case (cf. Subsection 5.3), hence $BL0+BL1$ ($CL0+CL$) is inconsistent.
@@
  • Consistency of BL0. The expression $\mu(x,y)$ is (indeed) a binary operation in BL, so that new operations $\to$ and $\leftrightarrow$ are well-defined. Therefore, BL0 is an extension by definition, hence a conservative extension of BL, so that BL0 is consistent (relative to BL).
@@
$\vdash_{BL0}IRLA,$
i.e. BL0 is a super-system of IRLA.
@@
    Proof of $x=y\Rightarrow x\leftrightarrow y=1$ and more. By definition:
@@
    If $x\neq y$, then $x\nleq y$ or $y\nleq x$, so $x\to y=0$ or $y\to x=0$, thus, let’s say $y\to x=0$, $x\leftrightarrow y=(x\to y)\wedge(y\to x)=(x\to y)\wedge 0=0$.
@@
    Proof of $x\to y=x\wedge y\leftrightarrow x$.
    If $x\leq y$ i.e. $x\wedge y=x$ by definition of $\leq$, then $x\to y=1$ and $x\wedge y\leftrightarrow x=1$, so it holds.
    If $x\nleq y$ i.e. $x\wedge y\neq x$, then $x\to y=0$ and $x\wedge y\leftrightarrow x=0$, so it also holds.
@@
    Proof of $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$.
@@
$\vdash_{BL1}IRLA,$
i.e. BL1 is a super-system of IRLA.
@@
    Proof of $x=y\Rightarrow x\leftrightarrow y=1$. If $x=y$, then $x\leftrightarrow y=x\leftrightarrow x=(x\to x)\wedge(x\to x)=x\to x=\neg x\vee x=1$.
@@
    Proof of $x\to y=x\wedge y\leftrightarrow x$. We have $x\wedge y\leftrightarrow x=(x\wedge y\to x)\wedge(x\to x\wedge y))=(\neg(x\wedge y)\vee x))\wedge(\neg x\vee(x\wedge y)=(\neg x\vee\neg y\vee x)\wedge(\neg x\vee(x\wedge y))=1\wedge(\neg x\vee y)=\neg x\vee y=x\to y$.
@@
    Proof sketch of $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$.

### v12

@@ 10 $\vDash\chi[\phi\mapsto\psi]$ 9: Definition
Suppose a conditional proof $\phi\vdash\psi$ of IRL satisfies that if a member of its proof sequence is a theorem then it is a tautology, then $\phi\to\psi$ is a tautology. That is, the rule of conditional proof (the CP rule) is tautologousness-preserving.
@@
Only the three primitive rules need considered in the meta proof: Conjunction introduction $\{\phi,\psi\}\Rightarrow\phi\wedge\psi$, Conjunction elimination $\phi\wedge\psi\Rightarrow\phi$, and Replacement property $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$.
@@
    • Case 1: $\psi$ is a theorem. In this case, $\psi$ is a tautology by the “theorem-is-tautology” condition, so $\phi\to\psi$ must be a tautology as shown below:
@@ 7     $\vDash\phi\to\chi$ 6: Definition
    9     $\vDash\phi\to\chi$ 8: Induction hypothesis (as $\vDash\phi\to\psi_{k}$ for all $k=1,...,n$)
@@ 10 $\vDash\phi\to\chi$ 2,3–7,8–9: Proof by exhaustion
    This result is used in the following four cases for $\psi$.
@@ 11 $\vDash\phi\to\psi$ 10: Definition
    Thus, for all $n\geq 1$, suppose that for all $k=1,...,n$, any conditional proof $\phi\vdash\psi$ of length $k$ satisfying the “theorem-is-tautology” condition results in tautological $\phi\to\psi$, then any length-$(n+1)$ conditional proof $\phi\vdash\psi$ satisfying the “theorem-is-tautology” condition leads to tautological $\phi\to\psi$ as well.
@@
From Meta Theorem 4.4.1, Meta Theorem 4.4.2 and Meta Theorem 4.4.3, it can be concluded that IRL is sound with respect to its semantics specified by IRLA.
@@
    All axioms of IRLA that are inherited from Boolean lattice have their counterparts in IRL. For example, IRLA axiom $x\wedge y=y\wedge x$ corresponds to IRL axiom $\phi\wedge\phi\leftrightarrow\psi\wedge\phi$.
@@
    The IRLA axiom $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$ (one of the two conditions for defining $\leftrightarrow$ in IRLA) corresponds to IRL theorem $(\phi\leftrightarrow\psi)\wedge\chi\to\chi[\phi\mapsto\psi]$ (from the primitive rule $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$ of IRL).
@@
    The IRLA axiom $x\to y=x\wedge y\leftrightarrow x$ (the defining axiom for $\to$) corresponds to IRL axiom $(\phi\to\psi)\leftrightarrow(\phi\wedge\psi\leftrightarrow\phi)$.
@@
For all the two primitive rules of IRLA, namely $x=y\Rightarrow x\leftrightarrow y=1$ and $\{s=t,u=v\}\Rightarrow(u=v)[s\mapsto t]$, if their premises have corresponding theorems in IRL, so are their conclusions.
@@
    $x=y\Rightarrow x\leftrightarrow y=1$: In IRL, $\phi\leftrightarrow\psi$ is a theorem if and only if $(\phi\leftrightarrow\psi)\leftrightarrow\top$ is a theorem by the rule of theorem representation. Thus, if the premise $x=y$ has a corresponding theorem $\phi\leftrightarrow\psi$ in IRL, then the conclusion $x\leftrightarrow y=1$ has the counterpart theorem $(\phi\leftrightarrow\psi)\leftrightarrow\top$ in IRL.
@@
    $\{s=t,u=v\}\Rightarrow(u=v)[s\mapsto t]$: For the premises $s=t$ and $u=v$, if their counterparts $E\leftrightarrow F$ and $G\leftrightarrow H$ are theorems of IRL, then $(G\leftrightarrow H)[E\mapsto F]$, the counterpart in IRL for the conclusion $(u=v)[s\mapsto t]$, must be a theorem in IRL as shown below:
    1 $\vdash E\leftrightarrow F$ Known
    2 $\vdash G\leftrightarrow H$ Known
    3     $E\leftrightarrow F$ 1 (assumed theorem)
    4     $G\leftrightarrow H$ 2 (assumed theorem)
    5     $(G\leftrightarrow H)[E\mapsto F]$ 3,4: Replacement (IRL rule)
    6 $\vdash(G\leftrightarrow H)[E\mapsto F]$ 3–5: Unconditional proof
From Meta Theorem 4.4.4 and Meta Theorem 4.4.5, it can be concluded that IRL is complete with respect to its semantics specified by IRLA.
@@ 4.5 Isomorphism between IRL and IRLA
Isomorphism between two systems can be equivalently expressed as that they are interpretable in each other in terms of the concept of interpretability (Tarski, 1953).
@@
Since IRL is sound and complete with respect to IRLA, there is a natural isomorphism between them as shown in Table 4.
@@
Formula schema $F(\phi,\psi,...)$ $t(x,y,...)$ Term
@@
As IRL is sound and complete, for a formula $\phi$ of IRL, “$\phi$ is valid or is a tautology” is equivalent to “$\phi$ is a theorem”, “$\phi$ is unsatisfiable or is a contradiction” is equivalent to “$\neg\phi$ is a theorem”; so “$\phi$ is a contingency” is equivalent to “both $\phi$ and $\neg\phi$ are not theorems” i.e. “$\phi$ is independent of IRL”.
@@ 4.6 Truth values of implication
Define $x<y$ as $x\leq y$ and $x\neq y$. Thus $10<x<$ 1 means that $x$ is (represents) a contingency.
By Corollary 4.3.2, in any cases we have the bounds of implication as in
$0\leq x\to y\leq\neg x\vee y,$
and both the lower bound 00 and the upper bound $\neg x\vee y$ of $x\to y$ are sharp, i.e. achievable, since we have $1\to 0=0$ and $x\to x=\neg x\vee x$.
@@
Thus, generally $V_{\phi\to\psi}$ is a subset of $V_{\neg\phi\vee\psi}$ (including of course $\emptyset$ and $V_{\neg\phi\vee\psi}$). Especially in the case where $\neg\phi\vee\psi$ is contingent, unlike classical logic that always adopts the equality $V_{\phi\to\psi}=V_{\neg\phi\vee\psi}$, $V_{\phi\to\psi}$ can be any subset of $V_{\neg\phi\vee\psi}$, i.e. $V_{\phi\to\psi}\in 2^{V_{\neg\phi\vee\psi}}$, so there are $2^{|V_{\neg\phi\vee\psi}|}$ subsets, of which only one is “correct” for $\phi\to\psi$. The correct $V_{\phi\to\psi}$ is determined by the original meaning of implication – a “certain mechanism” (cf. Subsection 2.2) corresponding to the implication $\phi\to\psi$ considered.
@@
Let $A$ and $B$ are primitive propositions having 4 possible truth-value combinations, i.e. the set of interpretations $U=\{\mathrm{TT,TF,FT,FF}\}$. Then $\emptyset\subset V_{\neg A\vee B}=\{\mathrm{TT,FT,FF}\}\subset U$, so $\neg A\vee B$ is contingent, therefore $V_{A\to B}\in 2^{V_{\neg A\vee B}}=2^{\{\mathrm{TT,FT,FF}\}}$, thus there are $82^{|\{\mathrm{TT,FT,FF}\}|}=2^{3}=$ 8 possibilities and only one of them is correct for $V_{A\to B}$.
The following is an example that shows which should be the “correct one” for the T-set of implication $\phi\to\psi$ as a subset of the T-set of $\neg\phi\vee\psi$.
@@ 7       F       F       F       T       F
    T-set of $\neg P\vee Q$ is the set $\{1,2,3,4,6,7\}$ of rows, indicating that it is contingent.
@@
    T-set of $P\to Q$, determined by T-set of $O$ rather than that of $P$ and $Q$, is the set $\{1,2,3\}$ of rows, a proper subset of that of $\neg P\vee Q$.
In this example, it is the certain mechanism of “in the open air” that corresponds to the inevitable implication relation of “if it rains, then the ground gets wet”, and the scope of the mechanism is the T-set of the implication. This is a typical example showing that it is not uncommon that $\neg\phi\vee\psi\Rightarrow\phi\to\psi$ is invalid. The example can explain why the classical logic, with the common “implication” unreasonably defined by $\phi\to\psi\Leftrightarrow\neg\phi\vee\psi$, is doomed to fail frequently in common situations in real world.
@@ 5.1 Definitions and notations
In the following, view a proof system as the set of its axioms and primitive rules, and use “statement” to mean formula or rule expression of a system; treat an object and the singleton containing the object the same thing.
@@
Definition 5.1.3 (System extension).
@@
An extension of $A$ with $X$, denoted as $A+X$, is the system obtained from $A$ by adding each member of $X$ as an axiom or a primitive rule.
@@
Note that $X\vdash_{A}Y$ is different from $\vdash_{A+X}Y$: the former is just the conditional proof of $Y$ from $X$ in $A$, while the latter is not a conditional proof in $A$ – at most we may define the meta conditional proof $(\vdash_{A}X)\vdash_{A}(\vdash_{A}Y)$ as $\vdash_{A+X}Y$.
@@
In terms of the above definitions: an extension of a system is a super-system of that system; the (deductively) equivalent relation “$=$” between two systems can be used in a usual way (i.e. it has the properties of reflexivity, symmetry, transitivity and replacement) ; the extension operator “$+$” is commutative and associative. Let $A$, $B$ and $C$ be sets of statements, any of which can be a system, then:
@@ 5.2 Two super-systems of IRLA
Let $BL=(X,1,0,\neg,\wedge,\vee,\leq)$ be the Boolean lattice described in Subsection 4.2. $BL0=(X,1,0,\neg,\wedge,\vee,\to,\leftrightarrow,\leq)$ and $BL1=(X,1,0,\neg,\wedge,\vee,\to,\leftrightarrow,\leq)$ are extensions of $BL$ defined as
@@
  • BL0 is isomorphic to a “contra-classical” logic, denoted as CL0, obtained by translating BL0 to a logic in terms of Table 4, so $BL0\cong CL0$. The logic CL0 is contra-classical because the definitions of implication $x\to y=\mu(x,y)$ in BL0 and $x\to y=\neg x\vee y$ in BL1 are contradictory in the general case (cf. Subsection 5.3), hence $BL0+BL1$ ($CL0+CL$) is inconsistent.
@@
  • Consistency of BL0. For any given $(x,y)\in X^{2}$, $\mu(x,y)\in X$ uniquely exists (either $1$ or 00), that is, $(x,y)\mapsto\mu(x,y)$ is (indeed) a binary operation on $X$, thus the new operations $\to$ and $\leftrightarrow$ are well-defined. Therefore, BL0 is an extension of BL by definition, hence a conservative extension of BL, thus BL0 is consistent relative to BL (so they are equiconsistent as BL0 is a super-system of BL).
@@
$\vdash_{BL0}IRLA.$
@@
    Proof of $x=y\Rightarrow x\leftrightarrow y=1$ (and $x\neq y\Rightarrow x\leftrightarrow y=0$) in BL0. By definition of $\to$ and $\leftrightarrow$ in BL0:
@@
    If $x\neq y$, then $x\nleq y$ or $y\nleq x$, so $x\to y=0$ or $y\to x=0$, thus, let’s say $x\to y=0$, $x\leftrightarrow y=(x\to y)\wedge(y\to x)=0\wedge(x\to y)=0$.
@@
    Proof of $x\to y=x\wedge y\leftrightarrow x$ in BL0.
    If $x\leq y$ i.e. $x\wedge y=x$ by definition of $\leq$, then $x\to y=1$ and $x\wedge y\leftrightarrow x=1$, so $x\to y=x\wedge y\leftrightarrow x$.
    If $x\nleq y$ i.e. $x\wedge y\neq x$, then $x\to y=0$ and $x\wedge y\leftrightarrow x=0$, so $x\to y=x\wedge y\leftrightarrow x$.
@@
    Proof of $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$ in BL0.
@@
$\vdash_{BL1}IRLA.$
@@
    Proof of $x=y\Rightarrow x\leftrightarrow y=1$ in BL1. If $x=y$, then $x\leftrightarrow y=x\leftrightarrow x=(x\to x)\wedge(x\to x)=x\to x=\neg x\vee x=1$.
@@
    Proof of $x\to y=x\wedge y\leftrightarrow x$ in BL1. We have $x\wedge y\leftrightarrow x=(x\wedge y\to x)\wedge(x\to x\wedge y)=(\neg(x\wedge y)\vee x))\wedge(\neg x\vee(x\wedge y))=(\neg x\vee\neg y\vee x)\wedge(\neg x\vee(x\wedge y))=1\wedge(\neg x\vee y)=\neg x\vee y=x\to y$.
@@
    Proof sketch of $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$ in BL1.